# OVERLAY 2025: 7th Workshop on Artificial Intelligence and Formal Verification, Logic, Automata, and Synthesis, October 26th, Bologna, Italy

### v1

@@ 1 Introduction
Recurrent Neural Networks (RNNs) are a class of Deep Neural Networks conceived primarily to process sequential data. Their linear complexity in space and time, combined with their efficiency, leads to their adoption in various applications from Time Series Forecasting [1] to policy modelling in reinforcement learning environments [2]. RNNs have been extensively studied from the lens of automata and formal language theory to model and analyze their behavior [3, 4]. Concretely, automata learning from RNNs seeks to map (or approximate) its semantics to some kind of automaton.
Automata learning methods are either based on active learning, where a form of feedback is received from the System Under Learning (SUL), which is the RNN in our case, or passive learning, where we only require a set of execution traces. Our work focuses on the latter approach. Our goal is to construct a symbolic automaton; more precisely, a Lattice Automaton, from a set of traces of an RNN. Lattice Automata are similar to Deterministic Finite Automata (DFA), however, they can recognize languages over an infinite alphabet, such as the inputs of an RNN, which are usually in the form of floating-point vectors. Therefore, we can be closer to its real behavior, and give a surrogate model of an RNN trained to recognize a series of real-world inputs (e.g. language modelling tasks such as sentiment analysis [5]).
@@
In recent years, there has been a renewed interest in learning automata from RNNs, for interpretability as well as for verification purposes. Active learning techniques are mainly based on extensions of the L* algorithm for learning regular languages. Multiple works have been proposed [6, 7, 8] to extend L* for extracting automata from RNNs.
On the other hand, passive learning techniques rely on the hypothesis that in RNNs, semantically similar intermediate neurons (also called hidden states) tend to cluster together [9]. Multiple techniques relying on clustering were proposed for learning DFA [10, 11], Weighted Finite Automata (WFA) [12], Hidden Markov Models (HMM) [13], and Probabilistic Finite Automata (PFA) [14]. The type of automata influences the complexity of the solution and is mainly motivated by the application; for instance, DFAs are well-suited to deterministic RNN behaviors, whereas WFAs and PFA are better alternatives to capture their probabilistic dynamics. A different approach by [15] relies on the classical Gold algorithm [16] by creating a prefix tree automaton (PTA) from the execution traces, then merging its states to yield a DFA. The authors associate RNNs hidden states values to each state of the PTA and then use that during the merge algorithm by only merging states that have at most a distance $d$ between their associated hidden states. In this paper, we adapt the latter method to learn automata over an infinite alphabet.
@@ 2 Key Notions
Any set $\Omega\subseteq\mathds{R}^{d}$ can be abstracted by a box ($d$-uple of intervals) $\alpha(\Omega)=\langle\omega_{1}\dots\omega_{d}\rangle$, such that each $\omega_{i}$ is an interval bounding the projection of $\Omega$ on the dimension $i$. It is a classical abstract domain, and there is a Galois connection [17] between $({\mathcal{P}}(\mathds{R}^{d}),\subseteq)$ and $(\Lambda,\sqsubseteq)$. This lattice is atomistic: its atoms are the singletons $\langle\llbracket x_{1},x_{1}\rrbracket\dots\llbracket x_{d},x_{d}\rrbracket\rangle$ for any vector $\langle x_{1}\dots x_{d}\rangle\in\mathds{R}^{d}$, thus the isomorphism between $\mathrm{Atoms}(\Lambda)$ and $\mathds{R}^{d}$. Because of this isomorphism, we identify any $x\in\mathds{R}^{d}$ with $\alpha(x)\in\Lambda$. That is why, in the definition of the language recognized by a Lattice Automaton, we wrote the condition $x_{i}\sqsubseteq\lambda_{i}$", while strictly speaking, the condition should be $\alpha(x_{i})\sqsubseteq\lambda_{i}$" since $x_{i}\in\mathds{R}^{d}$.
@@
A finite partition on $\mathrm{Atoms}(\Lambda)$ is a function $\Pi:\{1\dots n\}\rightarrow{\mathcal{P}}(\mathrm{Atoms}(\Lambda))$ such that $\bigcup_{i=1}^{n}\Pi(i)=\mathrm{Atoms}(\Lambda)$ and $i\neq j\Rightarrow\Pi(i)\cap\Pi(j)=\emptyset$. Since $\Lambda$ is atomistic, for any $i\in\{1\dots n\}$, we can identify the set of atoms $\Pi(i)$ with $\sqcup\{a|a\in\Pi(i)\}$ which defines the "maximal element" (also noted $\Pi(i)$ by abuse of notation) of this class. We have the property that classes of the partition are stable for the operation $\sqcup$ : $\forall\lambda_{1},\lambda_{2}:\lambda_{1}\sqsubseteq\Pi(i)\wedge\lambda_{2}\sqsubseteq\Pi(i)\Rightarrow\lambda_{1}\sqcup\lambda_{2}\sqsubseteq\Pi(i)$. Partitions are needed to properly define Lattice Automata, see [18] for details.
@@ 3 Learning Lattice Automata from RNN execution traces
The proposed algorithm is an extension of Gold’s algorithm for the case of ILA. We first create an Interval Prefix Tree Automaton (IPTA) from a set of traces of the RNN. Then, we launch the merging phase in which we merge states according to a similarity score inspired by [15].
@@ 4 Experiments
Our experiments were run on a Dell Inc. Precision 3591 computer, equipped with an Intel® Core™ Ultra 7 165H × 22 CPU. Python 3.12.3 was used for algorithms implementation and data synthesis tasks. We also used PyRAT [19] to define the abstractions.
We evaluated our algorithm on two benchmarks. First we run experiments with RNNs trained on Tomita languages, to compare our findings with [15] and demonstrate that our approach for inferring ILA is on par with SoTA. The fidelity to the original RNN of ILA are close to DFA but a bit less efficient (Figure 4), and since ILA are more complex than DFA, we have a larger number of states in our ILA. In that sense, our merging approach might benefit from further improvements to reduce the number of states. Then, we propose a novel benchmark by extending the Tomita languages to accept sequences of floats as inputs. For example, the first language accepts only sequences of numbers between 0 and $10$. The complete definition is given in Table 1.
@@ 4 No substring containing 3 consecutive letters $\in\llbracket-10,0\llbracket$
ILA size 9 5 23 5 7 17 82
ILA Fidelity 100 75.08 100 100 100 100 81.83
Type I error 0 0 0 0 0 0 16.05
Type II error 0 24.92 0 0 0 0 2.12
@@
RNN accuracy 99.98 98.64 99.94 98.95 99.89 95.91 96.35
ILA size 10 7 16 43 11 9 40
ILA Fidelity 82.33 62.75 99.14 86.86 88.07 82.75 71.53
Type I error 0.03 0.06 0.01 7.72 0.5 0.02 0
Type II error 17.64 37.19 0.85 5.42 11.43 17.23 28.47
@@
The scores of extracted ILA from a Tomita 2.0 RNN are roughly 80%, which is acceptable at this stage. The distribution of error scores suggests that our interval abstraction is precise enough (except for language 4, where it can reach 7.72%); however, type II error reflects an insufficient sample size, especially for language 2. Increasing the size of our sequences $S$ (which was set to 1000 sequences of maximum length 20 for each) will likely lead to improving the scores. Our preliminary results can be improved by further benchmarking the merge parameter and analyzing errors. Also by experimenting with different state merging algorithms that perform better than Gold’s algorithm, such as RPNI or EDSM [20], would lead to better results.
@@ 5 Conclusion and Future Work
We presented in this paper a passive learning algorithm that is capable of inferring an automaton from a set of traces of an RNN. Unlike previous methods, we do not require the inputs of the RNN to belong to a finite alphabet. Our algorithm ensures that we obtain an overapproximation of the set of traces. Our experiments demonstrate a capacity to infer ILA from an RNN trained on possibly infinite regular language, while also indicating that there remains significant potential for further improvements (increase fidelity and reduce ILA size), and to extend our benchmarks.

### v2

@@ 1 Introduction
Recurrent Neural Networks (RNNs) are a class of Deep Neural Networks conceived primarily to process sequential data. Their linear complexity in space and time, combined with their efficiency, leads to their adoption in various applications from Time Series Forecasting [lin2023segrnnsegmentrecurrentneural] to policy modelling in reinforcement learning environments [Schrittwieser_2020]. RNNs have been extensively studied from the lens of automata and formal language theory to model and analyze their behavior [Jacobsson2005RuleEF, Bollig2022ASO]. Concretely, automata learning from RNNs seeks to map (or approximate) its semantics to some kind of automaton.
Automata learning methods are either based on active learning, where a form of feedback is received from the System Under Learning (SUL), which is the RNN in our case, or passive learning, where we only require a set of execution traces. Our work focuses on the latter approach. Our goal is to construct a symbolic automaton; more precisely, a Lattice Automaton, from a set of traces of an RNN. Lattice Automata are similar to Deterministic Finite Automata (DFA), however, they can recognize languages over an infinite alphabet, such as the inputs of an RNN, which are usually in the form of floating-point vectors. Therefore, we can be closer to its real behavior, and give a surrogate model of an RNN trained to recognize a series of real-world inputs (e.g. language modelling tasks such as sentiment analysis [Johnson2016SupervisedAS]).
@@
In recent years, there has been a renewed interest in learning automata from RNNs, for interpretability as well as for verification purposes. Active learning techniques are mainly based on extensions of the L* algorithm for learning regular languages. Multiple works have been proposed [ANGLUIN198787, Mayr2021PropertyCW, Mukardin2022LearningFS] to extend L* for extracting automata from RNNs.
On the other hand, passive learning techniques rely on the hypothesis that in RNNs, semantically similar intermediate neurons (also called hidden states) tend to cluster together [muškardin2023relationshiprnnhiddenstate]. Multiple techniques relying on clustering were proposed for learning DFA [adax, Wang2018ACS], Weighted Finite Automata (WFA) [Wei2023WeightedAE], Hidden Markov Models (HMM) [Song2023LUNAAM], and Probabilistic Finite Automata (PFA) [Dong2019AnalyzingRN]. The type of automata influences the complexity of the solution and is mainly motivated by the application; for instance, DFAs are well-suited to deterministic RNN behaviors, whereas WFAs and PFA are better alternatives to capture their probabilistic dynamics. A different approach by [merrill2022extractingfiniteautomatarnns] relies on the classical Gold algorithm [GOLD1967447] by creating a prefix tree automaton (PTA) from the execution traces, then merging its states to yield a DFA. The authors associate RNNs hidden states values to each state of the PTA and then use that during the merge algorithm by only merging states that have at most a distance $d$ between their associated hidden states. In this paper, we adapt the latter method to learn automata over an infinite alphabet.
@@ 2 Key Notions
Any set $\Omega\subseteq\mathds{R}^{d}$ can be abstracted by a box ($d$-uple of intervals) $\alpha(\Omega)=\langle\omega_{1}\dots\omega_{d}\rangle$, such that each $\omega_{i}$ is an interval bounding the projection of $\Omega$ on the dimension $i$. It is a classical abstract domain, and there is a Galois connection [CC77] between $({\mathcal{P}}(\mathds{R}^{d}),\subseteq)$ and $(\Lambda,\sqsubseteq)$. This lattice is atomistic: its atoms are the singletons $\langle\llbracket x_{1},x_{1}\rrbracket\dots\llbracket x_{d},x_{d}\rrbracket\rangle$ for any vector $\langle x_{1}\dots x_{d}\rangle\in\mathds{R}^{d}$, thus the isomorphism between $\mathrm{Atoms}(\Lambda)$ and $\mathds{R}^{d}$. Because of this isomorphism, we identify any $x\in\mathds{R}^{d}$ with $\alpha(x)\in\Lambda$. That is why, in the definition of the language recognized by a Lattice Automaton, we wrote the condition $x_{i}\sqsubseteq\lambda_{i}$", while strictly speaking, the condition should be $\alpha(x_{i})\sqsubseteq\lambda_{i}$" since $x_{i}\in\mathds{R}^{d}$.
@@
A finite partition on $\mathrm{Atoms}(\Lambda)$ is a function $\Pi:\{1\dots n\}\rightarrow{\mathcal{P}}(\mathrm{Atoms}(\Lambda))$ such that $\bigcup_{i=1}^{n}\Pi(i)=\mathrm{Atoms}(\Lambda)$ and $i\neq j\Rightarrow\Pi(i)\cap\Pi(j)=\emptyset$. Since $\Lambda$ is atomistic, for any $i\in\{1\dots n\}$, we can identify the set of atoms $\Pi(i)$ with $\sqcup\{a|a\in\Pi(i)\}$ which defines the "maximal element" (also noted $\Pi(i)$ by abuse of notation) of this class. We have the property that classes of the partition are stable for the operation $\sqcup$ : $\forall\lambda_{1},\lambda_{2}:\lambda_{1}\sqsubseteq\Pi(i)\wedge\lambda_{2}\sqsubseteq\Pi(i)\Rightarrow\lambda_{1}\sqcup\lambda_{2}\sqsubseteq\Pi(i)$. Partitions are needed to properly define Lattice Automata, see [Gall2007LatticeAA] for details.
@@ 3 Learning Lattice Automata from RNN execution traces
The proposed algorithm is an extension of Gold’s algorithm for the case of ILA. We first create an Interval Prefix Tree Automaton (IPTA) from a set of traces of the RNN. Then, we launch the merging phase in which we merge states according to a similarity score inspired by [merrill2022extractingfiniteautomatarnns].
@@ 4 Experiments
Our experiments were run on a Dell Inc. Precision 3591 computer, equipped with an Intel® Core™ Ultra 7 165H × 22 CPU. Python 3.12.3 was used for algorithms implementation and data synthesis tasks. We also used PyRAT [lemesle2024neural] to define the abstractions.
We evaluated our algorithm on two benchmarks. First we run experiments with RNNs trained on Tomita languages, to compare our findings with [merrill2022extractingfiniteautomatarnns] and demonstrate that our approach for inferring ILA is on par with SoTA [^2] We identified and corrected implementation errors in the lattice automaton merging and minimization functions, as well as in training the RNNs. The results reported here supersede those in the workshop version.. Then, we propose a novel benchmark by extending the Tomita languages to accept sequences of floats as inputs. For example, the first language accepts only sequences of numbers between 0 and $10$. The complete definition is given in Table 1.
@@ 4 No substring containing 3 consecutive letters $\in\llbracket-10,0\llbracket$
ILA size 1 2 4 3 4 3 4
ILA Fidelity 100 100 100 100 100 100 100
Type I error 0 0 0 0 0 0 0
Type II error 0 0 0 0 0 0 0
@@
RNN accuracy 100 100 100 100 100 100 100
ILA size 1 2 4 3 4 3 4
ILA Fidelity 100 100 100 100 100 100 100
Type I error 0 0 0 0 0 0 0
Type II error 0 0 0 0 0 0 0
@@
The scores of extracted ILA from a Tomita 2.0 RNN are similar to state of the art compared to Tomita languages. It is trivial to see that the automata topologies and number of states for Tomita 2.0 is exactly the same as for Tomita. Our preliminary results can be improved by applying them for RNNs trained on real world tasks experimenting with different state merging algorithms that perform better than Gold’s algorithm, such as RPNI or EDSM [Soubki2023BenchmarkingSA].
@@ 5 Conclusion and Future Work
We presented in this paper a passive learning algorithm that is capable of inferring an automaton from a set of traces of an RNN. Unlike previous methods, we do not require the inputs of the RNN to belong to a finite alphabet. Our algorithm ensures that we obtain an overapproximation of the set of traces. Our experiments demonstrate a capacity to efficiently infer ILA from an RNN trained on possibly infinite regular language.